# DataSculpt: Crafting Data Landscapes for LLM Post-Training through Multi-objective Partitioning

### v1

@@ Abstract.
The effectiveness of long-context modeling is important for Large Language Models (LLMs) in various applications. Despite their potential, LLMs’ efficacy in processing long context does not consistently meet expectations, posing significant challenges for efficient management of prolonged sequences in training. This difficulty is compounded by the scarcity of comprehensive and diverse training datasets suitable for long sequences, which stems from inherent length biases across different data sources, and the logistical complexities associated with massive data management for training in extended contexts. In this work, we introduce DataSculpt, a data construction framework designed to strategically augment the data architecture for extended-context training. Our thorough evaluations demonstrate DataSculpt’s remarkable capacity to boost long-context training performance, achieving improvements including an 18.09% increase in retrieval augmentation, 21.23% in summarization, 21.27% in reading comprehension, and a 3.81% rise in code completion, all while preserving the models’ overall proficiency with a 4.88% improvement.
@@
Keer Lu, Zheng Liang, Xiaonan Nie, Da Pan, Shusen Zhang, Keshi Zhao, Weipeng Chen, Zenan Zhou, Guosheng Dong, Wentao Zhang, Bin Cui. PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX Corresponding Author.
@@
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@
The source code, data, and other artifacts will be released after the paper is accepted.
@@ 1. Introduction
Large Language Models (LLMs) have become a cornerstone in Artificial Intelligence (AI) (Brown et al., 2020; Achiam et al., 2023; Lewkowycz et al., 2022), particularly for Natural Language Processing (NLP) tasks, reshaping AI research and applications with unparalleled success across benchmarks and real-world scenarios (Dwivedi et al., 2021). Long context sequence enables LLMs to perform tasks significantly beyond existing paradigm (Fu et al., 2024), making it a pivotal area of research that has garnered substantial attention in recent years (Vaswani et al., 2017). However, LLMs’ efficacy in processing long context does not consistently meet the expected standards. Furthermore, despite claims of its feasibility, simply extending the training context window of LLMs does not guarantee an improvement in the long context modeling capabilities (Li et al., 2024). Hence, there is a compelling call for innovative strategies and methods to effectively cultivate the long context capabilities of LLMs.
Data plays a fundamental role during the training phases of LLMs (Miao et al., 2024; Fernandez et al., 2023; Whang et al., 2023). While the significance of model architectures is widely acknowledged, data-focused approaches hold substantial importance (Wang et al., 2023). The strategic management of data, especially in constructing an appropriately tailored training dataset, is critical for synergizing model performance and streamlining training processes across both pretraining and supervised fine-tuning stages (Wang et al., 2023; Kumar et al., 2017). Efforts in data management have been directed towards exploring ways to enhance model capabilities for processing long contexts. ICLM (Shi et al., 2023) situates related documents within the same context, thereby exposing LLMs to extensive relevant contexts and supplying training indicators that transcend document boundaries. Similarly, SPLICE (Staniszewski et al., 2023) utilized retrieval methodologies such as BM25 (Robertson et al., 2009) and Contriever-MSMARCO (Izacard et al., 2021) to execute structured packing within the training context, with a focus on code-related data. However, given that each document in the training corpus is accessed only once, the strategy employed in the previous approaches yields a notable quantity of isolated documents. As a result, during the later stages, this leads to the concatenation of unvisited documents with relatively low semantic correlation. Simultaneously, considering the current state of extended context sequences used for training, it leads to three key challenges:
Figure 1. Distribution of token counts for individual documents across different domains. Domains are arranged in descending order from top to bottom based on the proportion of documents with token numbers exceeding 64K. Documents within the domains of books and academic papers predominantly feature medium to long lengths (L $\geq$ 16K), in contrast to other fields, especially web-sourced data, where document lengths are almost exclusively below 4K.
C1. Insufficient Domain-Balanced Long Training Corpora. The challenge lies in effectively incorporating a wide range of domains into extended-context LLM training. Achieving a balanced domain mixture is imperative, since each domain encompasses distinct knowledge. However, the distribution of text lengths varies substantially across various domains, as depicted in Figure 1. While some domains offer data lengths suitable for long-context training, other domains, particularly web-sourced data where less than 1% of documents exceed 4K tokens, necessitate the concatenation of multiple documents to achieve adequate training contexts. This may lead to increased cross-attention among different documents within sequences, which is an undesirable learning signal.
C2. Absence of Criteria for Long Context Construction. While existing work has demonstrated the positive impact of structured data on amplifying language models’ long-context capabilities (Levine et al., 2021; Chan et al., 2022; de Vries, 2023a; Devlin et al., 2018), these efforts remain preliminary. Research dedicated to the organization and quality assessment of long-sequence training data for LLMs tends to focus on a narrow set of factors, lacking a systematic framework. Consequently, there is a discernible need for a detailed and cohesive framework that systematically organizes training data for applications requiring extensive context.
C3. Efficiency in Massive Data Management. Training LLMs with extended contexts from scratch is notably resource-intensive, requiring considerable time and labor investments (Chen et al., 2024a), which has redirected emphasis towards post-training optimizations as a more feasible strategy. Furthermore, organizing extensive pretraining datasets into long-context sequences poses substantial logistical hurdles (To et al., 2018), further compounding the complexity of developing LLMs efficiently. Consequently, devising and implementing strategies for efficient data scheduling and management is essential to enhance the capabilities of LLMs in a sustainable and cost-effective manner (Polyzotis et al., 2017), balancing the need for extensive context processing with the practical limitations of resource availability.
To address these issues, we introduce DataSculpt, a novel framework designed to reconfigure the organization of extended-context training data into a multi-objective combinatorial optimization problem, as illustrated in Figure 2. DataSculpt meticulously balances several critical objectives throughout the data organization process, including relevance, homogeneity, integrity, as well as computational efficiency, adopting a coarse-to-fine methodology to optimize the architecture of training data. We conducted post-training with 15 billion tokens on a 7B model across varying sequence lengths of 16K, 32K, and 64K, and experimental results demonstrate that DataSculpt improves performance in an array of tasks, most significantly in general understanding (+4.88%), retrieval augmentation (+18.09%), summarization (+21.23%), reading comprehension (+21.27%) and code completion (+3.81%). The finding reveals that DataSculpt can markedly synergize the model’s performance in long-context utilization during the post-train phase, even with a constrained number of training tokens.
Our contributions can be summarized as follows:
  • We highlight the challenge of compiling an appropriate long-context training corpus with a balanced domain mixture, crucial for enhancing the capabilities of LLMs in processing extensive text sequences.
  • To create suitable long context training data, we propose an efficient framework DataSculpt, which employs a multi-objective partition allocation strategy for the construction of extended context sequences.
  • Through quantitative analysis and t-SNE visualization of clustering outcomes at the intermediary stage of DataSculpt, we validate DataSculpt’s robustness for multi-objective data concatenation within extended sequences.
  • We engaged in continual pretraining using DataSculpt with various context lengths on a 7B model, and subsequently assessed its performance across 17 distinct tasks, illuminating the profound impact of our proposed method.
2. Related Work
2.1. Training Long-Context Language Models
Training Large Language Models (LLMs) with expanded context windows facilitates the execution of tasks that surpass the capabilities defined by the existing paradigm, such as multi-document question answering (Caciularu et al., 2023), long document summarization (Bai et al., 2023), and enhanced retrieval functionalities (Liu et al., 2024). Nevertheless, the endeavor of training LLMs with extended contexts is notably time-consuming and labor-intensive. Furthermore, documents with long sequences are conspicuously scarce in the training corpus (de Vries, 2023a). For instance, less than 5% of documents within CommonCrawl exceed 2k tokens in length (Cafarella and Halevy, 2011), which hinders LLMs’ exposure to complex narratives. Despite these challenges, significant strides have been made toward enhancing LLMs’ capabilities for long context processing.
Architecture-Based Strategies.  Numerous studies explored the variants of Transformer architecture by modifying its attention mechanism, which includes introducing recurrence (Dai et al., 2019), transforming attention into less computationally demanding forms (Beltagy et al., 2020; Zaheer et al., 2020), and employing low-rank approximations (Wang et al., 2020; Peng et al., 2021). Furthermore, LONGLLAMAs (Tworkowski et al., 2024) presented the Focused Transformer (FOT) featuring a novel training objective that enhances the (key, value) structure, which facilitates the expansion of context length and enables efficient scaling through k-nearest neighbors lookup. These approaches achieve more efficient exact attention through the development of an IO-optimized CUDA kernel specifically designed for this purpose.
Data-Centric Approaches.  The effectiveness of long-context language models relies equally upon the innovative employment of data-centric strategies throughout their training processes. The construction of long-context data involves more than a mere concatenation of short documents, which can lead to cross-document attention within a sequence. In these circumstances, formulating high-quality or domain-specific long sequences holds significant research value (de Vries, 2023b). ICLM (Shi et al., 2023) and SPLICE (Staniszewski et al., 2023) combine similar training instances through a sophisticated retrieval process, which emphasizes the critical role of high-quality long-context training data in enhancing the context expansion capabilities of LLMs without necessitating extensive data resources. Additionally, Krell et al. (Krell et al., 2021) deliberate on the challenges associated with the conventional concat-and-chunk approach, proposing an approximate bin-packing algorithm as a potential solution. As an improvement, dataset decomposition (Pouransari et al., 2024) and Best-fit Packing (Ding et al., 2024) both propose the sequence length aware optimization technique to minimize truncations, thereby facilitating more efficient pretraining on long sequences and achieving scalability relative to the size of dataset. Moreover, ProLong (Chen et al., 2024a) introduces an evaluation framework for assessing the quality of long-context data, further enriching the field’s methodological toolkit. Nonetheless, these methods have not performed a systematic framework tailored for data construction in long context training. Our work conceptualizes data management for extended-context training as a multi-objective combinatorial optimization problem, and provide a comprehensive framework for long sequence data organization.
2.2. Multi-Objective Combinatorial Optimization (MOCO)
Multi-objective combinatorial optimization (MOCO) problems are characterized by the need to optimize multiple objectives simultaneously, which are often in conflict with one another (Ulungu and Teghem, 1994). The complexity of MOCO challenges lies in the fact that improving the outcome of one objective often leads to the deterioration of another, making it necessary to find a balance or trade-off between them.
MOCO-Augmented LLMs.  Existing works have integrated MOCO problems with relevant concerns in the LLMs field. Wang et al. (Wang et al., 2024) employed the MOCO approach within the Reinforcement Learning from Human Feedback (RLHF) framework for LLMs, where the reward models aggregate multiple reward objectives with coefficients into a scalar score. Akiba et al. (Akiba et al., 2024) viewed model merging as a MOCO challenge, and introduced an innovative approach that leverages evolutionary algorithms to facilitate the automated development of robust foundation models. Furthermore, integration of MOCO with data-centric LLMs have been explored for training data selection (Du et al., 2023; Xu et al., 2023; Chen et al., 2024a), which applies heuristic techniques to assign weights to different data evaluation metrics and facilitates a phased resolution approach. Considering the extensive scale of training dataset, data construction for long context is treated as a MOCO challenge, for which the coarse-to-fine algorithm is identified as a suitable resolution strategy.
3. Method
We propose an innovative approach to construct concatenated sequences with multiple attributes for the enhancement of LLMs long-context utilization.
3.1. Multi-Objective Partition Allocation
Figure 2. Illustration of DataSculpt. In the data preprocessing stage, documents are initially divided into chunks with $\leq$ context length $L$, which are subsequently transformed into vector embeddings. During the semantic clustering phase, we implement a variant of the ISODATA algorithm, which is augmented with the FAISS vector searching library, to aggregate documents based on semantic similarity. Following this, a greedy semantic-driven largest-fit algorithm is utilized to arrange these clusters into sequences that are multi-objective optimally configured for extended-context training.
We consider the problem as a multi-objective partition allocation problem with soft constraints in a massive data scale, which is essentially an instance of the multi-objective combinatorial optimization (MOCO) problem (Lin et al., 2022). This can be formulated as

### v2

# DataSculpt: Crafting Data Landscapes for Long-Context LLMs through Multi-Objective Partitioning

@@ Abstract.
In recent years, foundation models, particularly large language models (LLMs), have demonstrated significant improvements across a variety of tasks. One of their most important features is long-context capability, which enables them to process and generate extended text while maintaining coherence, retrieving relevant information, and efficiently handling tasks with substantial amounts of text. The key to improving long-context performance lies in effective data organization and management strategies that integrate data from multiple domains and optimize the context window during training. Through extensive experimental analysis, we identified three key challenges in designing effective data management strategies that enable the model to achieve long-context capability without sacrificing performance in other tasks: (1) a shortage of long documents across multiple domains, (2) effective construction of context windows, and (3) efficient organization of large-scale datasets.
To address these challenges, we introduce DataSculpt, a novel data management framework designed for long-context training. We first formulate the organization of training data as a multi-objective combinatorial optimization problem, focusing on attributes including relevance, homogeneity, integrity, and efficiency. Specifically, our approach utilizes a coarse-to-fine methodology to optimize training data organization both efficiently and effectively. We begin by clustering the data based on semantic similarity (coarse), followed by a multi-objective greedy search within each cluster to score and concatenate documents into various context windows (fine). Our comprehensive evaluations demonstrate that DataSculpt significantly enhances long-context training performance, resulting in improvements of $18.09\%$ in retrieval augmentation, $21.23\%$ in summarization, $21.27\%$ in reading comprehension, and a $3.81\%$ increase in code completion, while also maintaining overall model proficiency with a $4.88\%$ improvement.
@@
Keer Lu, Xiaonan Nie, Zheng Liang, Da Pan, Shusen Zhang, Keshi Zhao, Weipeng Chen, Zenan Zhou, Guosheng Dong, Bin Cui, Wentao Zhang. PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX Equal Contribution.
Corresponding Author.
@@
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/8023looker/DataSculpt.
@@ 1. Introduction
In recent years, artificial intelligence has achieved remarkable milestones across a wide range of fields (Dwivedi et al., 2021; Lewkowycz et al., 2022), including computer vision (Dosovitskiy, 2020; Liu et al., 2021; Riquelme et al., 2021), natural language processing (Brown et al., 2020; Devlin et al., 2018; Radford et al., 2019; Raffel et al., 2020) and generative AI (Achiam et al., 2023). These breakthroughs are the result of significant progress in the development of AI systems, which is built upon three critical components: data, algorithms, and computational resources. Data forms the foundation, serving as the essential input from which models extract and generalize patterns (Wang et al., 2023; Kumar et al., 2017). The quality, quantity, and diversity of the data are pivotal in shaping a model’s ability to perform effectively across a range of tasks. Algorithms define how data is processed and guide the architectural design of models, enabling them to address specific problems. Computational resources, such as GPUs, are vital for handling the complex computations required throughout the training process, ensuring scalability and efficiency as models grow in complexity (Nie et al., 2023; Miao et al., 2024).
Historically, advancements in AI were primarily driven by model-centric approaches, where the focus was on refining and optimizing algorithmic architectures, such as AlexNet (Krizhevsky et al., 2012), ResNet (He et al., 2016), GoogleNet (Szegedy et al., 2015) in convolutional neural networks. However, in recent years, there has been a paradigm shift from the model-centric approach to the data-centric approach for foundation models, such large language models (LLMs). This shift emphasizes the quality, organization, and management of data as the critical factor for improving model performance (Kumar et al., 2017; Wang et al., 2023; Whang et al., 2023). For example, state-of-the-art LLM models, such as ChatGPT (OpenAI, 2022), PaLM (Anil et al., 2023), Gemini (Team et al., 2023) and Mixtral (Jiang et al., 2023) all leverage similar transformer architectures, but they employ distinct data organization and management strategies. Therefore, effective training data organization and management has emerged as a decisive factor in scaling model training and enhancing model capabilities (Miao et al., 2024; Fernandez et al., 2023; Prakash et al., 2024; Whang et al., 2023), particularly as models continue to grow in complexity and size.
The long-context capability is one of the most important features in LLMs (Fu et al., 2024; Xiong et al., 2023; Li et al., 2023), referring to their ability to process and generate extended text while maintaining coherence, retrieving relevant information (Liu et al., 2024), and handling tasks that involve substantial amounts of text efficiently (Bai et al., 2023). For instance, GPT-4 (Achiam et al., 2023) can handle context windows of up to 128K tokens, and maintain thematic and logical consistency across long-form text. Achieving this long-context capability requires training the model on an extended context window, which usually consist of multiple documents of varying lengths and different domains.
The primary challenges lie in determining effective data organization and management strategies for long-context model training, which ensure the model achieves long-context capability without compromising its performance in other tasks, such as multitask language understanding (Huang et al., 2019). Specifically, the strategy of selecting appropriate documents from vast sources of varying lengths to compose each context window during training remains under-explored, yet it is crucial for maximizing model efficiency and performance. Existing methods primarily focus on the combination of document sources, aiming to enhance the model’s general understanding and generation capabilities. However, they often overlook the consideration of document length, limiting the model’s ability to effectively handle long-sequence modeling. For example, ICLM (Shi et al., 2023) groups related documents within the same context window, enabling LLMs to access a broader range of relevant contexts, while SPLICE (Staniszewski et al., 2023) employs retrieval-based strategies (Robertson et al., 2009; Izacard et al., 2021) to implement structured packing within the context window. Unlike previous approaches, we try to explore the data management strategy for long-context model training from a multi-objective combinatorial optimization problem, which involves both document source and length:
How to design a data organization and management strategy for LLMs that could achieve long-context capability without compromising the models’ performance in other tasks?
Through extensive experimental analysis, we identified that the challenges associated with data organization and management strategies stem from the following three key aspects:
C1: Shortage of Long Documents Across Multiple Domains. Achieving a balanced domain mixture in training data is essential, as each domain contributes distinct knowledge to the model. However, data from various domains exhibit significant variation in length distribution, complicating the creation of effective long-context training datasets. While some domains provide sufficiently long documents suitable for long-context training, others, particularly web-sourced data, contain shorter documents, with less than 1% exceeding 4K tokens, as shown in Figure 2. To create adequate training contexts from such data, it is often necessary to concatenate multiple documents. This process, however, introduces potential challenges, such as increased cross-attention between unrelated documents, leading to undesirable learning signals and higher computational costs. Consequently, the variation in document length across domains can hinder the model’s ability to fully leverage long-context capabilities without introducing noise or redundancy.
C2: Effective Construction for Context Window. Developing robust algorithms for constructing context windows is a critical challenge in enhancing the long-context capabilities of language models. These algorithms must intelligently group related documents to preserve semantic coherence while maximizing the utility of the available data. Poorly designed methods can result in the combination of unrelated or loosely connected documents, which diminishes the quality of training signals and ultimately leads to suboptimal model performance. Although existing research has shown the positive impact of structured data on improving language models’ long-context capabilities (Levine et al., 2021; Chan et al., 2022; de Vries, 2023; Devlin et al., 2018), these efforts remain in their infancy. Previous approaches (Shi et al., 2023) have frequently resulted in isolated documents, as each document in the training corpus was accessed only once. This limitation necessitates the concatenation of unrelated documents with low semantic correlation in subsequent stages, as illustrated inTable 2 and Figure 4(b). Moreover, these methods often neglect the essential task of preserving individual document integrity, concentrating on a narrow set of factors rather than adopting a comprehensive framework for context window construction.
C3: Efficient Organization for Large-Scale Datasets. Efficiently organizing large-scale datasets into long-context sequences presents a significant challenge (To et al., 2018; Chen et al., 2024a). This task involves not only managing vast amounts of data but also ensuring that the organizational process minimizes computational overhead and labor costs. Inefficient data management strategies can lead to increased training times, elevated computational expenses, and difficulties in effectively scaling the model. Consequently, optimizing the data organization process is crucial for developing more efficient and scalable large language models (Polyzotis et al., 2017; Pan et al., 2024; Abadi et al., 2009). Such strategies must ensure a balance between long-context processing and the practical constraints of resource availability, ultimately leading to improved performance and scalability.
To address these challenges, we introduce DataSculpt, a novel data management framework designed for long-context training, which formulates the organization of training data as a multi-objective combinatorial optimization (MOCO) problem. DataSculpt aims to strike a balance among several key objectives, including maximizing relevance and homogeneity among documents within the same context window (C1, C2), preserving document integrity when truncating text that exceeds the context window’s capacity (C2), and optimizing computational efficiency in handling large-scale datasets (C3). Specifically, our approach employs a coarse-to-fine methodology to optimize the organization of training data efficiently and effectively. We start by clustering the data based on semantic similarity (coarse), then perform a multi-objective greedy search within each cluster to score and concatenate documents into different context windows (fine).
To verify the effectiveness of our methods, we conducted continual pretraining with 15 billion tokens on a language model with 7 billion parameters, utilizing varying sequence lengths of 16K, 32K, and 64K. This approach enables the model to continuously adapt and refine its knowledge, resulting in enhanced performance over a one-time pretraining method, making it the most widely used. Our experimental results demonstrate that our DataSculpt significantly enhances performance across a range of tasks, including general understanding ($+4.88\%$), retrieval augmentation ($+18.09\%$), summarization ($+21.23\%$), reading comprehension ($+21.27\%$), and code completion ($+3.81\%$). These findings indicate that DataSculpt effectively synergizes the model’s performance in long-context utilization during the continual pretraining phase, even with a limited number of training tokens.
Paper Organization. In Section 2, we introduce the background concepts relevant to this work, including transformer training and the multi-objective combinatorial optimization problem. Then, we present a quantitative analysis of the training datasets in Section 3, emphasizing the challenges faced in long-context LLM training. In Section 4, we introduce DataSculpt, which uses a coarse-to-fine methodology for efficiently organizing training data, enabling the model to achieve long-context capability without compromising performance on other tasks. Finally, we validate the effectiveness of DataSculpt through comprehensive experiments in Section 5.
2. Preliminary
2.1. Transformer Architecture
Transformers are designed to address sequence modeling and transduction tasks, such as language modeling and machine translation (Vaswani et al., 2017). At the heart of their architecture lie self-attention networks and point-wise feed-forward networks, which are integral to each layer. These models perform dense algebraic operations, including matrix multiplications, which incurs significant computational and memory demands.
Self-Attention Network.  Self-attention stands for a critical component of the Transformer architecture, allowing the model to weigh the importance of different words in a sequence, irrespective of their positions (Zhang et al., 2019; Miao et al., 2022; Vaswani et al., 2017). For each word in a sequence, self-attention computes a score that represents how much focus to place on other parts of the input sequence when encoding that specific word. This mechanism involves three sets of weights: query, key, and value, which are learned during training. Mathematically, self-attention can be described as a mapping of a query and a set of key-value pairs to an output, where the output is a weighted sum of the values. Each weight assigned to a value is computed by a compatibility function of the query with the corresponding key. This allows the Transformer to capture complex dependencies and relationships within the data, making it highly effective for tasks that require an understanding of the entire sequence, such as text translation or summarization.
Feed-Forward Network.  Each transformer layer includes a position-wise feed-forward network (FFN), which is applied to each position separately and identically. This consists of two linear transformations with a ReLU activation in between. The purpose of the FFN is to transform the representation independently at each position with the same, learned linear transformation. This means that while the self-attention layers share information across positions in the input, the feed-forward network processes each position independently. This component is crucial as it introduces non-linearity into the model, enabling it to learn more complex functions. Moreover, it ensures that the Transformer can adapt to different data types and tasks, providing additional flexibility and power to the model (Vaswani et al., 2017).
2.2. LLMs Training
Large Language Models (LLMs) denote Transformer-based architectures equipped with hundreds of billions (or more) of parameters (Zhao et al., 2023), initially introduced to address sequence modeling and transduction challenges. Below, we have outlined several key concepts about LLMs training.
Pretraining.  The pretraining phase of LLMs is a crucial period for developing their general knowledge capabilities. During this phase, LLMs undergo unsupervised training with large-scale raw text to cultivate their language modeling skills. Given a sequence $\mathrm{\textbf{x}}=\{x_{1},x_{2},...,x_{n}\}$, the primary training task for LLMs at this stage is to predict subsequent tokens $x_{i}$ based on preceding tokens $\mathrm{\textbf{x}}_{<i}$ as context, outlined in Fig. 1. The objective function aims to maximize the likelihood of Equation 1.
(1) $\mathcal{L}_{LLM}^{PT}(\mathrm{\textbf{x}})=\sum_{i=1}^{n}\log{P(x_{i}|\mathrm{\textbf{x}}_{<i})}$
Continual Pretraining.  Continual pretraining is a phase that extends the initial pretraining, using the same objective as shown in Equation 1. The key distinction is that continual pretraining builds on an already trained model rather than starting from scratch. At this stage, the LLM has basic language modeling capabilities; the aim now is to enhance its specialized functions, such as long context modeling, leveraging the foundational skills acquired during the initial pretraining phase. This phase is increasingly seen as a practical approach to improving model performance without the exorbitant computational costs associated with starting anew, as argued in recent studies (Ibrahim et al., 2024; Gupta et al., 2023; Parmar et al., 2024).
Context Window.  The context window is the span of text that a model actively considers when predicting the next word or deciphering the text’s structure. This window defines the range of contextual neighbors used in calculating vector representations (Lison and Kutuzov, 2017). During both the pretraining and continual pretraining phases, raw texts from various documents are concatenated to form a continuous sequence that matches the length of the context window. Any text exceeding this length is truncated. The size of the context window is pivotal, as it directly impacts the model’s capacity for language comprehension and generation.
Figure 1. An illustration of the data processing workflow, training task, context window and self-attention mechanism.
Figure 2. An illustration on the shortage of long documents across multiple domains. Left: The length distribution of individual documents across various domains, arranged in descending order by the proportion of documents exceeding 64K tokens. Documents, such as books and academic papers predominantly conist of medium to long lengths (L $\geq$ 16K), whereas other domains, particularly web-sourced data, contain documents almost exclusively shorter than 4K tokens. Right: The proportion of each domain in the overall training dataset.
Table 1. Token statistics across domains. We analyzed the maximum, minimum, and average token counts for different domains.
Min Len 667 322 127 173 696 289 132 23
Avg Len 176,156 216,969 20,555 2,578 6,216 886 358 224
2.3. Multi-Objective Combinatorial Optimization
The multi-objective combinatorial optimization (MOCO) problem involves optimizing several objectives simultaneously, which typically conflict with each other (Ulungu and Teghem, 1994) and can be formulated as: